### v1

Abstract
Remote sensing image (RSI) interpretation typically faces challenges due to the scarcity of labeled data, which limits the performance of RSI interpretation tasks. To tackle this challenge, we propose EarthSynth, a diffusion-based generative foundation model that enables synthesizing multi-category, cross-satellite labeled Earth observation for downstream RSI interpretation tasks. To the best of our knowledge, EarthSynth is the first to explore multi-task generation for remote sensing. EarthSynth, trained on the EarthSynth-180K dataset, employs the Counterfactual Composition training strategy to improve training data diversity and enhance category control. Furthermore, a rule-based method of R-Filter is proposed to filter more informative synthetic data for downstream tasks. We evaluate our EarthSynth on scene classification, object detection, and semantic segmentation in open-world scenarios, offering a practical solution for advancing RSI interpretation.
Figure 1: A diffusion-based generative foundation model, EarthSynth, pretrained on multi-source and multi-category data, synthesizing Earth observation with a semantic mask and text for downstream remote sensing image interpretation tasks.
1 Introduction
Remote sensing image (RSI) interpretation sun2021research is fundamentally constrained by challenges such as severe class imbalance sharma2025addressing and limited availability of high-quality labeled data, significantly impeding the development of robust models for downstream tasks. However, labeling RSIs typically requires domain-specific expertise and substantial manual effort, making large-scale annotation time-consuming and costly. Consequently, an important research objective is to effectively exploit existing labeled Earth observation (EO) datasets by uncovering latent relationships among samples to improve data efficiency.
In parallel, recent advances in generative data augmentation antoniou2017data (GDA), particularly diffusion models rombach2022high (DMs), offer a promising avenue for synthesizing high-quality labeled training data. GDA using DMs has been widely applied across various domains chen2024comprehensive . Prior works he2022synthetic ; trabucco2023effective investigate text-to-image (T2I) models to boost classification performance in data-scarce scenarios, enlightening and inspiring the remote sensing community. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. By augmenting rare classes and enriching data diversity, these models can play a critical role in mitigating data scarcity and enhancing the generalization capabilities of RSI interpretation.
In the remote sensing community, most existing GDA approaches are trained predominantly on single-source EO, which inherently limits the diversity of generated categories and lacks generality and flexibility across broader remote sensing applications. However, training generative models on multi-source data requires balancing the quality and diversity of generated data. Txt2Img-MHN xu2023txt2img attempts to use GANs salimans2016improved to generate satellite images. DiffusionSat khanna2023diffusionsat and CRS-Diff tang2024crs propose conditional DMs for generating optical RSIs, incorporating diverse texture-based conditions and applying the synthesized data to downstream tasks such as road extraction. GeoSynth toker2024satsynth joint data manifold of images and labels for satellite semantic segmentation. AeroGen tang2024aerogen and MMO-IG yang2025mmo try to use synthesized training data for satellite object detection. However, these methods are typically trained on single tasks or single-source data, requiring repeated training and generation to meet diverse needs, which hinders their direct application in real-world scenarios.
To solve the above problems, we propose EarthSynth, a diffusion-based generative foundation model, synthesizing EO with a semantic mask and text for downstream RSI interpretation tasks as shown in Figure 1. First, we construct the EarthSynth-180K with multi-source and multi-category data, 180K samples for training EarthSynth. Specifically, we collect samples from multiple datasets and apply random cropping and category-augmentation strategy to standardize image resolution, ensuring alignment among images, semantic masks, and text descriptions. To our knowledge, EarthSynth-180K is the first large-scale remote sensing dataset for diffusion training. During training, we adopt the Counterfactual Composition (CF-Comp) strategy to simultaneously enhance layout controllability and category diversity, thereby enabling the generation of more informative EO data. For training data synthesis, a rule-based method of R-Filter is proposed to filter more informative synthetic data. We evaluate our EarthSynth with multiple datasets on scene classification, object detection, and semantic segmentation. Furthermore, the effectiveness of our method is demonstrated through comprehensive ablation studies and visualization analysis.
We summarize the main contributions as follows:
  • We propose EarthSynth, a diffusion-based generative foundation model trained on the EarthSynth-180K dataset with 180K samples aligned image, semantic mask, and text, achieving a unified solution to achieve multi-task generation.
  • EarthSynth employs the CF-Comp strategy to balance the layout controllability and category diversity during training, enabling fine layout control for RSI generation. It further incorporates the R-Filter to extract more informative and high-quality synthesized samples.
  • EarthSynth is evaluated on scene classification, object detection, and semantic segmentation, three RSI interpretation tasks in open-world scenarios.
2 Related Work
Diffusion Models for Remote Sensing. Diffusion models, which have shown great success in natural image synthesis rombach2022high , are increasingly being applied to remote sensing khanna2023diffusionsat ; liu2025text2earth . Depending on the use case liu2024diffusion , their applications in remote sensing can be broadly divided into three categories: image enhancement, image interpretation, and image synthesis. For image enhancement, DMs improve multispectral and hyperspectral images by fusing information across channels and expanding attention xu2023dual ; feng2024multiscale ; wei2024diffusion ; li2023hyperspectral . Some works also apply DMs to tasks like change detection bandara2022ddpm and climate prediction gao2023prediff , enabling better integration of multimodal data and improving pixel-level discrimination. On the other hand, image synthesis methods focus on generating artificial data to address the issue of data scarcity. Recent studies tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth demonstrate how synthetic data can benefit downstream remote sensing tasks. Unlike prior approaches that fine-tune on a single-source dataset for one specific task, we adopt a more general training and synthesis strategy. Our approach supports multiple tasks—scene classification, object detection, and semantic segmentation—allowing the model to generate more diverse and widely applicable data.
Figure 2: EarthSynth is trained with CF-Comp training strategy on real and unrealistic logical mixed data distribution, learns remote sensing pixel-level properties in multiple dimensions, and builds a unified process for conditional diffusion training and synthesis.
Generative Data Augmentation with Diffusion Models. Generative data augmentation using DMs has been explored across various domains, including natural images and remote sensing. Studies such as he2022synthetic ; trabucco2023effective investigate using text-to-image DMs to boost classification performance under limited data conditions. To further improve generation quality, alemohammad2024self ; alemohammad2024selfconsuming propose leveraging self-generated data; however, this approach risks model collapse due to overfitting. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. In remote sensing, recent works tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth have begun to explore the potential of diffusion-based data synthesis and enhancement to improve RSI interpretation tasks, like object detection or segmentation. But there is no unified solution to achieve multi-task generation.
3 Preliminaries
Training Data Synthesis. Given a conditional generative diffusion model $G_{\theta}$ with pretrained parameters $\theta$, and $x$ denotes the generative image. The the generative distribution w.r.t. $x$ induced by the condition set $\mathcal{C}$ is defined as:
$\mathcal{X}_{\mathcal{C}}=\sum_{i=1}^{|\mathcal{C}|}\frac{1}{|\mathcal{C}|}\,\mathbb{P}_{G_{\theta}}(x\mid c_{i}),$ (1)
where $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$ consists of a conditional Mask-Text pairs $c_{i}$, and each conditional distribution $\mathbb{P}_{G_{\theta}}(x\mid c_{i})$ is induced from $x=G_{\theta}(\epsilon\mid c_{i})$ by sampling $\epsilon\sim\mathcal{N}(0,I)$. The pixel-level, region-level and semantic-level label $y=\{y_{l},y_{c}\}=\mathcal{F}(c_{i})$ is generated via a mapping function $\mathcal{F}$, where $y$ consists of a location label $y_{l}$ (semantic mask or bounding box) and a category label $y_{c}$.
Feature Decomposition. Feature decomposition gao2023out for satellite imagery $x=f\left(x_{\texttt{obj}},x_{\texttt{bg}},x_{\texttt{noise}}\right)$ across various categories from different satellites. In remote sensing, we can formalize some criteria by (in)dependence with label $y$ in the meta distribution $\mathbb{P}_{G_{\theta}}$:
[…]
where $y$ depends on object $x_{\texttt{obj}}$ and background $x_{\texttt{bg}}$ but is independent of noise $x_{\texttt{noise}}$ which yields $\mathbb{P}_{G_{\theta}}(y\mid x)=\mathbb{P}_{G_{\theta}}(x_{\texttt{obj}},x_{\texttt{bg}})$. $x_{\texttt{noise}}$ is the noise disturbance during satellite imaging that makes semantic confusion pan2023reducing ; pan2023prior , leading to inter-class similarity and significant intra-class variation of satellite images.
[…]
4 Counterfactual Composition for Controllable Diffusion Training
Due to the inherent characteristics of satellite imagery, satellite images often exhibit high inter-class similarity and significant intra-class variation, posing challenges for RSI interpretation. To get more informative data distribution from a generative DM, we aim to approximate a real-world distribution $\mathcal{D}^{\texttt{real}}$ with as much training data distribution $\mathcal{D}^{\texttt{train }}\subset\mathcal{D}^{\texttt{real}}$ as possible. Constructing training data with diverse backgrounds and objects enables the DM to learn rich and diverse semantic information in open remote sensing scenarios. To achieve this, we enhance scene diversity through counterfactual composition, which involves combining existing object categories with diverse background contexts. The definition of counterfactual composition is as follows.
Figure 3: Copy-Paste operation of CF-Comp Strategy and Rule-based Filtering with CLIP.
Definition 1 (Counterfactual Composition).
Given a set of source elements ${A_{1},A_{2},\dots,A_{n}}$, where each $A_{i}$ represents a specific semantic component (e.g., object, region, or attribute) extracted from a distinct instance, a counterfactual sample is constructed by recombining these components as:
$x^{\prime}=\mathcal{CF}(A_{1}^{(i)},A_{2}^{(j)},\dots,A_{n}^{(k)}).\quad i\neq j\neq k$ (3)
Here, $A_{i}^{(i)}$ denotes the $i$-th component drawn from the $i$-th source instance, and $\mathcal{CF}(\cdot)$ is a composition function that logically combines elements from different instances to form a counterfactual sample $x^{\prime}$, where $x^{\prime}$ is out of the distribution of $\mathcal{D}^{\texttt{real}}$, being plausible yet not observed in reality.
Remark 1.
Counterfactual Composition allows for the generation of novel and logically consistent scenes by fusing diverse elements. The goal is to preserve semantic and structural coherence while expanding the diversity and improving generalization of possible inputs.
Unrealistic Logical Scene. We aim to construct an unrealistic logical scene by counterfactual composition, i.e., satellite imagery of real-world data that does not exist in the real world but is logically present. This also intended to maintain consistency in the image-label union space and prevent disruptions caused by arbitrary counterfactual composition. As shown in Figure 2, we use the Copy-Paste to perform counterfactual composition. For example, the unrealistic logical scene is obtained by combining the two images’ small and large vehicle objects in Figure 3. An unrealistic logical scene can be judged from three dimensions: channel, pixel, and semantic space. We define three criteria to measure whether two images can be combined into an unrealistic logical scene:
From the channel space dimension, the Image Color Sensitivity (ICS) can be obtained as
$\texttt{ICS}(x_{a},x_{b})=\begin{cases}\mathbf{1}_{\left\{\left|S(x^{a})-S(x^{b})\right|<s_{0}\right\}},&\text{if }C_{a}=C_{b}\\ 0,&\text{if }C_{a}\neq C_{b}\end{cases},$ (4)
where the color sensitivity $S=\operatorname{Var}(R-G)+\operatorname{Var}(R-B)+\operatorname{Var}(G-B)$, $C_{a}$ and $C_{b}$ are the number of channels of satellite images $x_{a}$ and $x_{b}$ respectively.
From the pixel space dimension, the Mask Overlap Rate (MOR) can be obtained as
$\texttt{MOR}(m^{a},m^{b})=\begin{cases}\dfrac{|m^{a}\cap m^{b}|}{|m^{a}\cup m^{b}|},&\text{if }|m^{a}\cup m^{b}|>0\\ 0,&\text{otherwise}\end{cases}.$ (5)
From the semantic space dimension, the text semantic similarity (TSS) can be obtained as
$\texttt{TSS}(t^{a},t^{b})=\frac{t^{a}\cdot{t^{b}}^{\top}}{\|t^{a}\|\cdot\|t^{b}\|}.$ (6)
Input: Training dataset $\mathcal{D}_{\texttt{train}}=\{(\mathcal{I},\mathcal{M},\mathcal{T})\}_{i=1}^{|\mathcal{D}_{\texttt{train}}|}$, where $\mathcal{I}$ is the ground-truth image, $\mathcal{M}$ is the mask prompt, and $\mathcal{T}$ is the text prompt
Output: Conditional diffusion model $\mathcal{M}_{\theta}$ with parameters $\theta$
for each training step do
       sample a mini-batch Image-Mask-Text triples $\mathcal{B}_{\texttt{ori}}=(\mathcal{I},\mathcal{M},\mathcal{T})$ from dataset $\mathcal{D}$;
      $\triangleright$ Counterfactual Composition
      $B_{\texttt{copy}}\leftarrow\{\}$, a thresholds $\alpha_{0}$, $\beta_{0}$, $\eta_{0}$
      for $(x^{a},m^{a},t^{a}),(x^{b},m^{b},t^{b})$ in $(\mathcal{I},\mathcal{M},\mathcal{T})\times(\mathcal{I},\mathcal{M},\mathcal{T})$  do
             $\alpha=\texttt{ICS}(x^{a},x^{b})$, $\beta=\texttt{MOR}(m^{a},m^{b})$, $\eta=\texttt{TSS}(t^{a},t^{b})$;
            if $\alpha$, $\beta$, $\eta$ > $\alpha_{0}$, $\beta_{0}$, $\eta_{0}$ then
                  $(x^{\prime},m^{\prime},t^{\prime})=\texttt{Copy-Paste}(x^{a}_{\texttt{obj}},x^{b})$;
                  $B_{\texttt{copy}}\leftarrow(x^{\prime},m^{\prime},t^{\prime})$;
      Get a new mini-batch Image-Mask-Text triples $\mathcal{B}=\mathcal{B}_{\texttt{ori}}+\mathcal{B}_{\texttt{copy}}$;
      Encode text prompt $\mathcal{T}$ using a frozen text encoder to obtain $E_{\mathcal{T}}$;
      Sample noise $\epsilon\sim\mathcal{N}(0,1)$ and timestep $t$;
      Generate noisy image: $x_{t}=\sqrt{\alpha_{t}}\cdot\mathcal{I}+\sqrt{1-\alpha_{t}}\cdot\epsilon$;
      Extract mask features $F_{\mathcal{M}}$ from $\mathcal{M}$ using $\mathcal{M}_{\theta}$ with parameters $\theta$;
      Inject mask features into a frozen UNet and predict noise $\hat{\epsilon}$:
           $\hat{\epsilon}=\text{UNet}(x_{t},t,E_{\mathcal{T}},\mathcal{M}_{\theta}(\mathcal{M}))$;
      Compute global loss and local loss: $\mathcal{L}=\|\hat{\epsilon}-\epsilon\|^{2}+\gamma\|F_{\mathcal{M}}\cdot\hat{\epsilon}-F_{\mathcal{M}}\cdot\epsilon\|^{2}$;
      Update $\theta$ using gradient descent to minimize $\mathcal{L}$;
Algorithm 1 Controllable Diffusion Training with Counterfactual Composition
Mixed Data Distribution. In real remote sensing scenarios, the object and background of an image typically follow a consistent data distribution. However, in unrealistic logical data distributions, discrepancies between foreground and background distributions often lead to out-of-distribution hsu2020generalized . By integrating real-world and unrealistic logical data distributions, the latent data manifold becomes more complex and diverse, allowing the model to generalize better to unseen scenarios while preserving the essential characteristics of the original distribution. Assuming that the object and background of the image can be modeled as Gaussian distributions, respectively, as:
$x_{\texttt{obj}}\sim\mathcal{N}(\mu_{\texttt{obj}},\sigma^{2}_{\texttt{obj}}I),\quad x_{\texttt{bg}}\sim\mathcal{N}(\mu_{\texttt{bg}},\sigma^{2}_{\texttt{bg}}I),$ (7)
where $(\mu_{\texttt{obj}},\sigma^{2}_{\texttt{obj}})$ and $(\mu_{\texttt{bg}},\sigma^{2}_{\texttt{bg}})$ denote the respective mean and variance of the object and background distributions. Assuming independence and linearity of expectation, the mean and variance of the resulting image $x^{\prime}$ with mask $m$ by Copy-Paste can be derived as:
$\mu_{\texttt{mix}}$ $=\alpha\mu_{\texttt{obj}}+(1-\alpha)\mu_{\texttt{bg}}$ (8)
$\sigma^{2}_{\texttt{mix}}$ $=\alpha\sigma^{2}_{\texttt{obj}}+(1-\alpha)\sigma^{2}_{\texttt{bg}}+\alpha(1-\alpha)(\mu_{\texttt{obj}}-\mu_{\texttt{bg}})^{2},$
where $\alpha=\frac{1}{HW}\sum_{i,j}m_{i,j}$ represents the proportion of foreground pixels in the image. The additional term $\alpha(1-\alpha)(\mu_{\texttt{obj}}-\mu_{\texttt{bg}})^{2}$ in the variance reflects the distributional shift caused by the mismatch between object and background statistics. CF-Comp allows models to learn richer and more informative representations by controlling the Copy-Paste ratio $\Sigma\alpha$ in remote sensing. We note that in satellite images, $\mu_{\texttt{obj}}\approx\mu_{\texttt{bg}}$ and $\sigma^{2}_{\texttt{obj}}\approx\sigma^{2}_{\texttt{bg}}$, but in natural scene images, these parameters may differ significantly, resulting in distributional shifts that the model must learn to accommodate.
Global and Local Loss. In contrast to binary masks, semantic masks embed category-specific information, making pixel-level precision essential. However, global constraints alone are inadequate for modeling such fine-grained control. Therefore, we design a diffusion-based training framework that integrates global and local loss constraints to achieve more accurate and semantically aligned generation. We extract semantic mask features $F_{\mathcal{M}}$ from mask $\mathcal{M}$ using condition model $\mathcal{M}_{\theta}$ with parameters $\theta$ and inject semantic mask features into a frozen UNet ronneberger2015u and predict noise $\hat{\epsilon}$:
$\mathcal{L_{\texttt{global}}}=\|\hat{\epsilon}-\epsilon\|^{2},\quad\mathcal{L_{\texttt{local}}}=\|F_{\mathcal{M}}\cdot\hat{\epsilon}-F_{\mathcal{M}}\cdot\epsilon\|^{2},\quad\mathcal{L}=\mathcal{L_{\texttt{global}}}+\gamma\mathcal{L_{\texttt{local}}}$ (9)
where $\gamma$ is the local constraint factor.
Training Process. Algorithm 1 shows the CF-Comp method for condition diffusion EarthSynth training. We follow Zhang et al. zhang2023adding to train a conditional diffusion model.
5 Training Data Synthesis
Since the quality of data generated by DMs can vary significantly, we propose a rule-based filtering method, R-Filter, to further refine the generated samples and retain only those that meet predefined quality criteria. This ensures that only high-quality and task-relevant samples are included in the final dataset.
Condition Sampling. During the data synthesis stage, conditions $c_{i}=(m_{i},t_{i})$ are randomly sampled by category from the training condition set $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$. We also apply random rotation, scaling and merging, based on the mask and text prompts, to get more diverse conditions.
Label Mapping. Different label mapping functions $\mathcal{F}$ are employed for different downstream tasks. For scene classification, the category label $y_{c}$ is directly obtained from the associated text $t_{i}$. For semantic segmentation, the semantic mask $m_{i}$ and the category label $y_{c}$ are derived by extracting class mappings from the corresponding mask. For object detection, bounding boxes are generated by extracting the contour features of the masks by using Ramer-Douglas-Peucker cao2022automatic algorithm.
Rule-based Filtering. We propose R-Filter, a rule-based method that uses CLIP scores to evaluate Image-Object-Background triplets $[x,x_{obj},x_{bg}]$ by computing overall, object-specific, and background scores, as shown in Figure 3. Since both $x_{obj}$ and $x_{bg}$ are related to the label $y$, we retain samples with high overall or object-specific scores for training downtown models with setting the CLIP score threshold $S_{0}$.
Synthesis Process. Algorithm 2 of the appendix shows training data synthesis with EarthSynth for scene classification, object detection, and semantic segmentation.
Table 1: Zero-shot scene classification accuracy on RSICD and DIOR datasets.
Method Scene Classification
RSICD* (Top1 / Top5) DIOR (Top1 / Top5)
Txt2Img-MHN(VQVAE) xu2023txt2img + CLIP 32.7 / 75.5 - / -
Txt2Img-MHN(VQGAN) xu2023txt2img + CLIP 40.9 / 72.7 - / -
CRS-Diff tang2024crs + CLIP 57.1 / 79.0 - / -
StableDiffusion rombach2022high + CLIP 61.3 / 88.3 41.5 / 73.0
InstanceDiffusion wang2024instancediffusion + CLIP 59.1 / 88.1 44.5 / 79.0
ControlNet zhang2023adding + CLIP 55.5 / 85.5 46.5 / 78.5
EarthSyth + CLIP 60.0 / 91.8 49.0 / 80.0
Table 2: Object detection on DOTAv2 and DIOR dataset with mAP.
Method Data Usage Object Detection
DOTAv2 DIOR
+ GroundingDINO Real 56.3 74.0
StableDiffusion rombach2022high + GroundingDINO Real + Synth - -
ControlNet zhang2023adding + GroundingDINO Real + Synth 57.4 (+1.1) 74.1 (+0.1)
EarthSyth + GroundingDINO Real + Synth 58.4 (+2.1) 74.3 (+0.3)
6 Experiment
In this section, we evaluate EarthSynth on scene classification, object detection, and semantic segmentation, including performance analysis, ablation studies, and visualization analysis.
EarthSynth-180K. EarthSynth-180K is built from OEM xia2023openearthmap , LoveDA wang2021loveda , DeepGlobe demir2018deepglobe , SAMRS wang2023samrs , and LAE-1M pan2025locate datasets, and enhanced with mask and text prompts. We leverage many pixel-level mask annotations and semantic-level texts as prompts for DMs. By applying Random Cropping Strategy and Category-Augmentation Strategy to the EarthSynth-180K dataset, we obtain about 180K high-quality Image-Mask-Text triplets consisting of images, semantic masks, and texts. This dataset has a wide range of categories and is labeled with semantic segmentation, which can be used to improve object detail reconstruction and category understanding in DMs. More details can be found in the Section A.1.
Table 3: Semantic segmentation on Potsdam, FloodNet, FAST, and FLAIR datasets with mIoU.
Method Data Usage Semantic Segmentation
Potsdam FloodNet FAST FLAIR
+ GSNet Real 40.6 33.9 16.8 19.3
StableDiffusion rombach2022high + GSNet Real + Synth - - - -
ControlNet zhang2023adding + GSNet Real + Synth 35.4 (-5.2) 39.3 (+5.4) 17.7 (+0.9) 20.4 (+1.1)
EarthSynth + GSNet Real + Synth 42.7 (+2.1) 44.4 (+10.5) 17.4 (+0.6) 21.6 (+2.3)
Evaluation. We use multiple tasks to evaluate data generation capability, including scene classification cheng2017remote , object detection zou2023object , and semantic segmentation guo2018review . We construct an evaluation task that progresses from coarse-grained to fine-grained levels, spanning from image-level to pixel-level, to assess the generalization capability of synthetic data.
Experiment Setup. All experiments are performed using four NVIDIA A100 GPUs, and the complete training of EarthSynth requires approximately 4 $\times$ 45 GPU hours. EarthSynth is initialized with the pretrained Stable Diffusion v1-5 rombach2022high weights. Training uses mixed precision to improve computational efficiency and reduce memory consumption. A batch size of 8 per device is used, with gradient accumulation over four steps, resulting in an adequate batch size of 32. In the CF-Comp setting, we set $s_{0}=150$, $\alpha_{0}=1$, $\beta_{0}=0.02$, and $\eta_{0}=0.6$. The local constraint $\gamma$ is set 10. The training runs for 40,000 steps. A constant learning rate of 1e-5 is adopted without any warm-up phase, and gradient clipping with a maximum norm of 1 is applied to ensure training stability. To improve data quality, we use the CLIP-ViT-B/32 model radford2021learning to filter out low-quality samples. The CLIP score threshold $S_{0}$ is set 0.4.
6.1 Comparative Results
We evaluate downstream remote sensing tasks in open-world scenarios, including scene classification, object detection, and semantic segmentation. Training data is synthesized using remote sensing-specific methods trained on RSICD lu2017exploring Txt2Img-MHN xu2023txt2img , CRS-Diff tang2024crs , and baseline methods trained on EarthSynth-180K, Stable Diffusion rombach2022high , InstanceDiffusion wang2024instancediffusion , ControlNet zhang2023adding . We adopt open-vocabulary methods, CLIP radford2021learning , GroundingDINO liu2024grounding , and GSNet ye2025towards for downtown evaluation method. Real denotes real-world data, and Synth refers to diffusion-generated data.
Scene Classification. We uniformly generated 10 images per category and averaged the results over three runs for evaluation. Figure 2 presents the comparative results of zero-shot scene classification regarding Top-1 and Top-5 accuracy on RSICD lu2017exploring and DIOR li2020object datasets. RSICD* is a subset of the original RSICD dataset containing 11 classes, used to align different dataset settings of different methods that do not include RSICD dataset. Compared to VAE-based and GAN-based methods, Diffusion-based approaches exhibit a clear advantage in generation quality. On RSICD, our method ranks second in Top-1 accuracy and first in Top-5 accuracy for classification. It also outperforms the baseline ControlNet by 6.3. Our method achieves superior scene classification performance on DIOR, indicating its ability to generate images with higher classification scores. These results suggest the potential of our approach for pretraining in downstream tasks.
Object Detection. We evaluate the effectiveness of our proposed method on object detection by fine-tuning the GroundingDINO model with 256 synthetic images per category. Table 2 experiments on the DOTAv2 xia2018dota and DIOR li2020object datasets show that our approach outperforms ControlNet, achieving improvements of 2.1 and 0.3 compared to training solely on real data. These results demonstrate the benefit of incorporating generated data for object detection. Few-shot detection results are shown in Section A.6, where our method consistently improves performance. Sample generated data can be found in Section A.7.
Semantic Segmentation. We generated 256 synthetic images for all categories of EarthSynth-180K for training the open-set segmentation method GSNet. Table 3 reports the mIoU results on four semantic segmentation datasets: Potsdam ISPRS2013 , FloodNet rahnemoonfar2021floodnet , FAST wang2023samrs , and FLAIR garioud2023flair . The baseline GSNet model achieves solid performance using only real data, particularly on the Potsdam dataset, and incorporating synthetic data from ControlNet results in mixed outcomes, with performance improvements on FloodNet by 5.4, FAST by 0.9, and FLAIR by 1.1. Still, a degradation of Potsdam by 5.2, suggesting potential issues related to multi-source data alignment. In contrast, EarthSynth consistently improves performance across all datasets, with gains of 2.1 on Potsdam, 10.5 on FloodNet, 0.6 on FAST, and 2.3 on FLAIR. These results demonstrate the effectiveness of EarthSynth-generated data in enhancing semantic segmentation under diverse scenarios.
Figure 4: Effect of training samples per class on the DOTAv2 dataset.
Table 4: Detection performance on DOTAv2 using different modules configurations.
Method mAP
*only real 56.3
baseline -
+ $\mathcal{L}_{\text{local}}$ 56.5
+ $\mathcal{L}_{\text{local}}$ + R-Filter 57.4
+ CF-Comp -
+ CF-Comp + $\mathcal{L}_{\text{local}}$ 57.9
+ CF-Comp + $\mathcal{L}_{\text{local}}$ + R-Filter 58.4
6.2 Ablation Studies
In this section, we peform the ablation studies of submodule, sample size and CF-Comp.
Submodule. We fix the random seed during inference to ensure consistency across comparative experiments, guaranteeing that identical templates are used throughout the evaluation. As shown in Table 4, object detection performance on DOTAv2 using different submodule configurations. We found that introducing the local loss $\mathcal{L}_{\texttt{local}}$ can learn layout control of sematic mask. Adding R-Filter further boosts performance to 57.4. Incorporating CF-Comp alongside $\mathcal{L}_{\texttt{local}}$ brings a larger gain, reaching 57.9. The full configuration with CF-Comp, $\mathcal{L}_{\text{local}}$, and R-Filter achieves the best result of 58.4 mAP. CF-Comp and R-Filter effectively identifies and retains higher-quality samples and synthetic more informative training data for downstream tasks.
Sample Size. Figure 4 shows that performance improves as the number of synthetic samples per class increases, reaching the highest mAP of 58.4 at 128 samples per class. However, further increases lead to marginal declines, suggesting that moderate-scale synthesis is most effective, while excessive data may introduce redundancy or noise.
CF-Comp. Further ablation studies on CF-Comp in the Section A.3.
6.3 Visualization Analysis
Synthesis. Figure 5 shows the synthesis satellite images with FID scores heusel2017gans on DOTAv2 dataset. Compared to the baseline ControlNet, our method results in larger distributional shifts and generates higher-quality images. In addition, by combining prompts with transformed masks, we can produce more robust images for training downstream tasks.
Embedding. Figure 6 shows the t-SNE van2008visualizing distributions and FID scores for three representative categories, illustrating the alignment between synthetic and real data. The use of CF-Comp increases sample diversity but also leads to larger distributional shifts, as indicated by higher FID scores. The introduction of R-Filter also increases this gap, with FID scores consistently increasing across three categories. These results suggest that CF-Comp and R-Filter enhances the semantic diversity and distributional spread of synthetic samples, potentially improving generalization for downstream tasks.
Figure 5: Visualization of synthesis satellite images on DOTAv2 dataset.
Figure 6: Embedding visualizations for some categories on the EarthSynth-180K dataset.
7 Discussion and Limitation
Training Cost. Counterfactual combinations do not take into account the time cost of actual training and are not applicable in cases where computational resources are scarce.
Generation efficiency. The generation process is executed using real multi-GPU inference, and the CLIP-based filtering introduces additional time cost of generation.
8 Conclusion
We propose EarthSynth, a diffusion-based generative foundation model trained on the EarthSynth-180K dataset. To balance layout controllability and category diversity during training, EarthSynth adopts the CF-Comp strategy, enabling precise layout control for RSI generation. Additionally, it incorporates the R-Filter post-processing method to extract more informative synthesized samples for downstream tasks, offering a practical and scalable solution to advance RSI interpretation through synthetic data.
Appendix A Technical Appendices and Supplementary Material
A.1 EarthSynth-180K
EarthSynth-180K is derived from OEM, LoveDA, DeepGlobe, SAMRS, and LAE-1M datasets. It is further enhanced with mask and text prompt conditions, making it suitable for training foundation DMs. The EarthSynth-180K dataset is constructed using the Random Cropping and Category-Augmentation strategies. The category distribution of the EarthSynth-180K dataset is presented in Table 7, along with the corresponding category-to-abbreviation mapping shown in Table 5.
Figure 7: Category distribution of the EarthSynth-180K dataset.
Table 5: Category to abbreviation mapping.
Category Abbr. Category Abbr. Category Abbr.
background BG ground track field GF vehicle VE
bare land BL small vehicle SV windmill WM
grass GR baseball diamond BD expressway service area EA
pavement PA tennis court TC expresswalltoll station ET
road RO roundabout RA dam DM
tree TR storage tank ST golf field GO
water WA harbor HA overpass OP
agriculture land AL container crane CC stadium SD
buildings BU airport AP train station TS
forest land FL helipad HP large vehicle LV
barren land BA chimney CH swimming pool SP
urban land UL helicopter HE bridge BR
plane PL ship SH soccer ball field SF
basketball court BC
Figure 8: A category-augmentation strategy to construct multiple object-background pairs.
Random Cropping Strategy. To standardize the input resolution for the DM, we employ a random cropping strategy to generate 512$\times$512 image patches. The same cropping operation is applied to the corresponding semantic masks to maintain spatial consistency. Specifically, if the image size is larger than the target crop size, a random top-left coordinate $(x,y)$ is selected to extract a patch of the desired dimensions. If the image is smaller than the crop size, no cropping is performed, and the region starting from the top-left corner is used directly. To construct textual descriptions for the categories, we employ a template-based approach by appending category names to the phrase “A satellite image of [class1], [class2], …”, resulting in descriptions such as “A satellite image of swimming pool, ship, small vehicle, harbor”. We construct around 50K triplets of images, textual descriptions, and semantic masks. Note that center-based cropping is not required, as the categories in the dataset are heterogeneous and spatially scattered.
Category-Augmentation Strategy. Specifically, to enhance the model’s understanding of individual categories and enable fine-grained category-specific control during generation, we perform data augmentation on each image to construct multiple foreground-background distinguishable pairs. As illustrated in Figure 8, these augmented triplets are generated from the previously obtained image-text-mask triplets by isolating non-background categories in the original masks and corresponding textual descriptions, resulting in a set of new, category-focused triplets.
A.2 Remote Sensing Diffusion Models
Table 6 compares remote sensing diffusion models across three prompt levels: semantic, region, and pixel. Most methods support only semantic-level prompts, offering global but coarse control. Some models introduce region-level prompts using bounding boxes to enhance spatial precision. Pixel-level prompts, such as masks and sketches, provide the most detailed control and are used by CRS-Diff, GeoSynth, and EarthSynth. EarthSynth uniquely combines semantic and pixel-level prompts, enabling both high-level semantics and fine-grained spatial guidance. This reflects a shift toward more precise and controllable image generation.
High-Resolution Synthesis. Text2Earth, DiffusionSat, and GeoSynth are representative methods aimed at generating high-resolution satellite imagery. These models leverage diffusion-based generative frameworks or text-to-image architectures to reconstruct fine-grained spatial details, often guided by auxiliary inputs such as text descriptions, semantic maps, or multi-modal signals. Their primary applications lie in image restoration, super-resolution, cloud removal, and spectral enhancement, which are essential for improving the visual and analytical quality of satellite data in scientific and operational settings.
Task-Oriented Synthesis. CRS-Diff, SatSynth, AeroGen, and MMO-IG are designed with downstream utility in mind, focusing on generating synthetic data tailored for specific tasks such as land cover classification, object detection, and change detection. These models incorporate task-specific priors, including class labels, semantic layouts, or instance-level masks, to guide the generation process. By aligning the synthesized data with the needs of target tasks, these methods enhance model generalization in low-resource scenarios, enable domain adaptation, and facilitate pretraining or fine-tuning of deep learning models in remote sensing applications.
Table 6: Comparison of remote sensing diffusion methods based on different levels of prompts.
Method Condition Type
Semantic-level (class, text) Region-level (box) Pixel-level (mask, sketch)
Text2Earth [17]
DiffusionSat [12]
GeoSynth [25]
CRS-Diff [13]
SatSynth [14]
AeroGen [15]
MMO-IG [16]
EarthSynth
A.3 EarthSynth Training Process Analysis
Local Loss. As shown in Figure 9 and Figure 10, the loss function is defined as
$\mathcal{L}=\mathcal{L}_{\texttt{global}}+\gamma\mathcal{L}_{\texttt{local}}.$
Setting $\gamma=10$ accelerates convergence and helps the model better capture semantic mask information. Compared to not using $\mathcal{L}_{\texttt{local}}$, incorporating it enables more effective layout control. We found it difficult to learn layout control without local constraints on mask. For comparison fairness, ControlNet+$\mathcal{L}_{\texttt{local}}$ is used in object detection and semantic segmentation.
CF-Comp. An ablation study on ICS and TSS further shows their complementary roles. Using ICS encourages the generation of more diverse and novel combinations, while using TSS favors more semantically reasonable compositions. It also validates the effectiveness of our EarthSynth approach using CF-Comp.
Figure 9: EarthSynth over time-step diffusion training process.
Figure 10: EarthSynth over time-step diffusion training process.
A.4 Training Data Synthesis Process
To construct a high-quality training dataset for remote sensing tasks, we adopt a conditional diffusion-based synthesis pipeline as described in Algorithm 2. Given a set of semantic and spatial conditions, our pre-trained diffusion model progressively generates synthetic samples from Gaussian noise through reverse diffusion steps. The process incorporates both text embeddings and mask features to guide generation. After sampling, we apply rule-based filtering using CLIP scores to retain only semantically meaningful and visually plausible samples. The resulting dataset combines diversity and control, providing effective data support for downstream training.
Input: Pre-trained conditional diffusion model $\mathcal{M}_{\theta}$ parameters $\theta$, sampling steps $T$, sampling condition set $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$
Output: Synthetic dataset $\mathcal{S}=\{x_{0}^{(i)},c_{i}\}_{i=1}^{|\mathcal{C}|}$
Initialize empty synthetic dataset: $\mathcal{S}\leftarrow\{\}$;
for each condition $c_{i}\in\mathcal{C}$ do
       Sample initial noise $x_{T}^{(i)}\sim\mathcal{N}(0,\mathbf{I})$;
      Encode text prompt $\mathcal{T}$ using a frozen text encoder to obtain $E_{\mathcal{T}}$;
      Extract mask features $F_{\mathcal{M}}$ from $\mathcal{M}$ using $\mathcal{M}_{\theta}$ with parameters $\theta$;
      for $t=T$ down to $1$ do
             Predict noise: $\hat{\epsilon}_{t}=\text{UNet}(x_{t}^{(i)},t,E_{\mathcal{T}},\mathcal{M}_{\theta}(c_{i}))$;
            Update latent variable using reverse diffusion equation:
                 $x_{t-1}^{(i)}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}^{(i)}-\sqrt{1-\alpha_{t}}\cdot\hat{\epsilon}_{t}\right)+\sigma_{t}z$,
                 where $z\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $z=0$;
      $\triangleright$ Rule-based Filtering
      if $Score_{\texttt{CLIP}}(x)>S_{0}\mid Score_{\texttt{CLIP}}(x_{obj})>S_{0}$ then
            Add generated sample to dataset: $\mathcal{S}\leftarrow\mathcal{S}\cup\{(x_{0}^{(i)},c_{i})\}$;
return Final dataset $\mathcal{S}$ sampled from conditional distribution $\mathcal{X}_{\mathcal{C}}$
Algorithm 2 Training Data Synthesis
A.5 Comparison of FID and CLIP scores
Table 7 compares fake data from different models using FID and CLIP scores. FID measures distribution distance to all images of the RSICD dataset, while CLIP evaluates semantic alignment. Results show that FID does not consistently align with CLIP, revealing different strengths in generation quality. This suggests that FID alone is not reliable for assessing task-specific data quality, as it lacks relevance to downstream performance.
Table 7: Comparison of FID and CLIP scores trained on single-source (Single) and multi-source Multi data. We calculated the FID score between the generated images and the RSICD dataset. * is from the original papers.
Type Method Training Dataset FID Score CLIP Score
Single StableDiffusion [4] RSICD 103.4 26.1
StableDiffusion [4] DIOR 228.1 26.2
InstanceDiffusion [34] RSICD 138.6 24.7
Text2Earth [17] Git-10M 24.5* -
DiffusionSat [12] fMoW 15.8* 17.2*
GeoSynth [25] US Cities 12.3* / 171.1 30.3* / 25.0
SatSynth [14] iSAID, LoveDA, OEM - -
MMO-IG [16] DIOR 34.5* -
AeroGen [15] HRSC,DIOR 38.6* -
Multi CRS-Diff [13] RSICD, fMoW 50.7* / 107.0 20.3* / 23.6
InstanceDiffusion [34] RSICD, RSITMD, UCMerced 123.1 25.0
ControlNet [32] EarthSynth-180K 183.8 25.9
EarthSynth EarthSynth-180K 198.7 26.1
A.6 Few-Shot Detection Results
Table 8 shows that open-source methods such as ETS [54] consistently outperform close-source baselines across all shot settings. For example, ETS achieves 37.5 percent mAP in the 10-shot setting, surpassing Detic at 15.4 percent and DE-ViT at 25.6 percent. Introducing synthetic data leads to notable improvements, especially when combined with real data. In the 5-shot setting, ETS improves from 29.3 percent using only real data to 33.9 percent with ControlNet-generated data and 34.1 percent with EarthSynth-generated data. EarthSynth consistently yields better results than ControlNet, suggesting higher-quality synthesis. These results highlight the effectiveness of augmenting real data with high-quality synthetic data for few-shot detection.
Table 8: The 1/5/10-shot detection results on the DIOR dataset.
Method Setting Data Usage mAP
1-shot Detic ViT-L [55] Close-Source Real 4.1
DE-ViT ViT-L [56] Close-Source Real 14.7
CD-ViTO [57] Close-Source Real 17.8
GroundingDINO Swin-B [45] Open-Source Real 11.7
ETS Swin-B [54] Open-Source Real 12.7
(ControlNet) ETS Swin-B [54] Open-Source Synth 9.2
(ControlNet) ETS Swin-B [54] Open-Source Real + Synth 13.1
(EarthSynth) ETS Swin-B [54] Open-Source Synth 9.5
(EarthSynth) ETS Swin-B [54] Open-Source Real + Synth 13.9
5-shot Detic ViT-L [55] Close-Source Real 12.1
DE-ViT ViT-L [56] Close-Source Real 23.4
CD-ViTO [57] Close-Source Real 26.9
GroundingDINO Swin-B [45] Open-Source Real 27.7
ETS Swin-B [54] Open-Source Real 29.3
(ControlNet) ETS Swin-B [54] Open-Source Synth 23.0
(ControlNet) ETS Swin-B [54] Open-Source Real + Synth 33.9
(EarthSynth) ETS Swin-B [54] Open-Source Synth 23.6
(EarthSynth) ETS Swin-B [54] Open-Source Real + Synth 34.1
10-shot Detic ViT-L [55] Close-Source Real 15.4
DE-ViT ViT-L [56] Close-Source Real 25.6
CD-ViTO [57] Close-Source Real 30.8
GroundingDINO Swin-B [45] Open-Source Real 36.4
ETS Swin-B [54] Open-Source Real 37.5
(ControlNet) ETS Swin-B [54] Open-Source Synth 25.3
(ControlNet) ETS Swin-B [54] Open-Source Real + Synth 40.2
(EarthSynth) ETS Swin-B [54] Open-Source Synth 26.4
(EarthSynth) ETS Swin-B [54] Open-Source Real + Synth 40.7
A.7 Examples of Generated Data
Figure 11 and Figure 12 present examples of the generated semantic segmentation and object detection data produced by the proposed method. From a visual perspective, segmentation labels are generally of higher quality compared to object detection labels, as the latter often rely on post-processing algorithms such as edge detection, which may introduce inaccuracies.
Figure 11: Some examples of generated semantic segmentation data.
Figure 12: Some examples of generated object detection data.
A.8 Comparison of Generated Images
CRS-Diff. Due to its limited training data, CRS-Diff produces lower-quality images compared to models trained on larger and more diverse datasets.
GeoSynth. GeoSynth is designed for high-resolution image generation based on data from US cities, but its capacity to generate a wide range of categories is limited.
Stable Diffusion. Stable Diffusion lacks the ability to generate layout-controllable images, which limits its applicability in tasks that require precise spatial arrangement of objects, such as semantic segmentation or object detection data synthesis.
ControlNet and EarthSynth. ControlNet and EarthSynth enable the generation of images with explicit layout control and visually salient targets, making them well-suited for structured data synthesis tasks.
Figure 13: Comparison of generated images with different methods.
A.9 Guidance Scale Analysis
Figure 14 shows how the guidance scale affects the CLIP score across two categories. As the scale increases, the CLIP Score generally rises at first, then levels off or slightly decreases. This suggests that moderate guidance scales lead to better alignment between the generated image and the text prompt. In the top examples, representing tennis courts, the CLIP Score reaches its peak around scale 4, with images showing clearer structure and improved object fidelity. In the bottom examples, representing playgrounds, the score is highest near scale 2 or 3, but have a poor image generated. Lower scales tend to produce blurry or semantically weak images, while higher scales enhance visual clarity but may reduce diversity. These results indicate that a moderate guidance scale, typically between 3 and 5, offers a good balance between semantic alignment and image quality. The results demonstrate that generation varies across different images, and tuning the guidance scale provides a simple way to control semantic accuracy and visual structure.
Figure 14: The guidance scale affects the CLIP score across two categories.