### v2

Abstract
Remote sensing image (RSI) interpretation typically faces challenges due to the scarcity of labeled data, which limits the performance of RSI interpretation tasks. To tackle this challenge, we propose EarthSynth, a diffusion-based generative foundation model that enables synthesizing multi-category, cross-satellite labeled Earth observation for downstream RSI interpretation tasks. To the best of our knowledge, EarthSynth is the first to explore multi-task generation for remote sensing, tackling the challenge of limited generalization in task-oriented synthesis for RSI interpretation. EarthSynth, trained on the EarthSynth-180K dataset, employs the Counterfactual Composition training strategy with a three-dimensional batch-sample selection mechanism to improve training data diversity and enhance category control. Furthermore, a rule-based method of R-Filter is proposed to filter more informative synthetic data for downstream tasks. We evaluate our EarthSynth on scene classification, object detection, and semantic segmentation in open-world scenarios. There are significant improvements in open-vocabulary understanding tasks, offering a practical solution for advancing RSI interpretation.
Figure 1: A diffusion-based generative foundation model, EarthSynth, pretrained on multi-source and multi-category data, synthesizing Earth observation with a semantic mask and text for downstream remote sensing image interpretation tasks.
1 Introduction
Remote sensing image (RSI) interpretation sun2021research is fundamentally constrained by challenges such as severe class imbalance sharma2025addressing and limited availability of high-quality labeled data, significantly impeding the development of robust models for downstream tasks. However, labeling RSIs typically requires domain-specific expertise and substantial manual effort, making large-scale annotation time-consuming and costly. Consequently, an important research objective is to effectively exploit existing labeled Earth observation (EO) datasets by uncovering latent relationships among samples to improve data efficiency.
In parallel, recent advances in generative data augmentation antoniou2017data , particularly Diffusion Models (DMs) rombach2022high , offer a promising avenue for synthesizing high-quality labeled training data. Data augmentation using DMs has been widely applied across various domains chen2024comprehensive . Prior works he2022synthetic ; trabucco2023effective investigate text-to-image (T2I) models huang2024learning to boost classification performance in data-scarce scenarios, enlightening and inspiring the remote sensing community. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. By augmenting rare classes and enriching data diversity NEURIPS2023_f99f7b22 , these models can play a critical role in mitigating data scarcity and enhancing the generalization capabilities of RSI interpretation. Unlike existing studies, our work centers on improving the generative diversity of diffusion models tailored for remote sensing applications.
In the remote sensing community, most existing data augmentation approaches are trained predominantly on single-source EO, which inherently limits the diversity of generated categories and lacks generality and flexibility across broader remote sensing applications. However, training generative models on multi-source data requires balancing the quality and diversity of generated data. Txt2Img-MHN xu2023txt2img attempts to use GANs salimans2016improved to generate satellite images. DiffusionSat khanna2023diffusionsat and CRS-Diff tang2024crs propose conditional DMs for generating optical RSIs, incorporating diverse texture-based conditions and applying the synthesized data to downstream tasks such as road extraction. GeoSynth toker2024satsynth joint data manifold of images and labels for satellite semantic segmentation. AeroGen tang2024aerogen and MMO-IG yang2025mmo try to use synthesized training data for satellite object detection. However, these methods are typically trained on single tasks or single-source data, requiring repeated training and generation to meet diverse needs, which hinders their application in real-world scenarios.
To solve the above problems, we propose EarthSynth, a diffusion-based generative foundation model, synthesizing EO with a semantic mask and text for downstream RSI interpretation tasks as shown in Figure 1. First, we construct the EarthSynth-180K with multi-source and multi-category data, 180K samples for training EarthSynth. Specifically, we collect samples from multiple datasets and apply random cropping and category-augmentation strategy to standardize image resolution, ensuring alignment among images, semantic masks, and text descriptions. To our knowledge, EarthSynth-180K is the first large-scale remote sensing dataset for diffusion training. During training, we adopt the Counterfactual Composition (CF-Comp) strategy with channel, pixel, and semantic spaces as batch-sample selection mechanism to simultaneously enhance layout controllability and category diversity, thereby enabling the generation of more informative EO data. Different previous studies, we apply the CF-Comp strategy to diffusion-based generative foundation model training for downtown tasks, avoiding repeating the training task-specific generative model for each downstream task. For training data synthesis, a rule-based method of R-Filter is proposed to filter more informative synthetic data. We evaluate our EarthSynth with multiple datasets on scene classification, object detection, and semantic segmentation. Furthermore, the effectiveness of our method is demonstrated through comprehensive ablation studies and visualization analysis.
We summarize the main contributions as follows:
  • We propose EarthSynth, a diffusion-based generative foundation model trained on the EarthSynth-180K dataset with 180K cross-satellite and multi-sensor samples aligned image, semantic mask, and text, achieving a unified solution to achieve multi-task generation.
  • EarthSynth employs the CF-Comp strategy to balance the layout controllability and category diversity during training, enabling fine layout control for RSI generation. And integrates the R-Filter post-processing method to extract more informative synthesized data.
  • EarthSynth is evaluated on scene classification, object detection, and semantic segmentation in open-vocabulary scenarios, well validating its effectiveness.
2 Related Work
Diffusion Models for Remote Sensing. Diffusion Models (DMs), which have shown great success in natural image synthesis rombach2022high , are increasingly being applied to remote sensing khanna2023diffusionsat ; liu2025text2earth . Depending on the use case liu2024diffusion , their applications in remote sensing can be broadly divided into three categories: image enhancement, image interpretation, and image synthesis. For image enhancement, DMs improve multispectral and hyperspectral images by fusing information across channels and expanding attention xu2023dual ; feng2024multiscale ; wei2024diffusion ; li2023hyperspectral . Some works also apply DMs to tasks like change detection bandara2022ddpm and climate prediction gao2023prediff , enabling better integration of multimodal data and improving pixel-level discrimination. On the other hand, image synthesis methods focus on generating artificial data to address the issue of data scarcity. Recent studies tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth demonstrate how synthetic data can benefit downstream remote sensing tasks. Unlike prior approaches that fine-tune on a single-source dataset for one specific task, we adopt a more general training and synthesis strategy. Notably, these data augmentation methods zhang2024advancing ; zhang2025controllable ; liu2025control ; li2025domain demonstrate significant advantages in few-shot learning tasks fu2023styleadv ; fu2025ntire . Our approach supports multiple tasks, including scene classification, object detection, and semantic segmentation, allowing the model to generate more diverse and widely applicable data.
Figure 2: EarthSynth is trained with CF-Comp training strategy on real and unrealistic data distribution, learns remote sensing pixel-level properties in multiple dimensions, and builds a unified process for conditional diffusion training and synthesis.
Generative Data Augmentation with Diffusion Models. Generative data augmentation using DMs has been explored across various domains, including natural images and remote sensing. Studies such as he2022synthetic ; trabucco2023effective investigate using text-to-image DMs to boost classification performance under limited data conditions. To further improve generation quality, alemohammad2024self ; alemohammad2024selfconsuming propose leveraging self-generated data; however, this approach risks model collapse due to overfitting. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. In remote sensing, recent works tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth ; li2025domain have begun to explore the potential of diffusion-based data synthesis and enhancement to improve RSI interpretation tasks. But there is no unified solution to achieve multi-task generation.
3 Preliminaries
Training Data Synthesis. Given a conditional generative diffusion model $G_{\theta}$ with pretrained parameters $\theta$, and $x$ denotes the generative image. The the generative distribution w.r.t. $x$ induced by the condition set $\mathcal{C}$ is defined as:
$\mathcal{X}_{\mathcal{C}}=\sum_{i=1}^{|\mathcal{C}|}\frac{1}{|\mathcal{C}|}\,\mathbb{P}_{G_{\theta}}(x\mid c_{i}),$ (1)
where $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$ consists of a conditional Mask-Text pairs $c_{i}$, and each conditional distribution $\mathbb{P}_{G_{\theta}}(x\mid c_{i})$ is defined by generating samples $x=G_{\theta}(\epsilon\mid c_{i})$, where $\epsilon\sim\mathcal{N}(0,I)$. The pixel-level, region-level, and semantic-level label $y=\{y_{l},y_{c}\}=\mathcal{F}(c_{i})$ is generated via a mapping function $\mathcal{F}$, where $y$ consists of a location label $y_{l}$ (semantic mask or bounding box) and a category label $y_{c}$.
Feature Decomposition. Feature decomposition gao2023out for satellite imagery $x=f\left(x_{\texttt{obj}},x_{\texttt{bg}},x_{\texttt{noise}}\right)$ across various categories from different satellites. In remote sensing, we can formalize some criteria by (in)dependence with label $y$ in the meta distribution $\mathbb{P}_{G_{\theta}}$:
$x_{\texttt{obj}},x_{\texttt{bg}}\not\!\perp\!\!\!\perp y,\quad x_{\texttt{noise}}\perp\!\!\!\perp y,\\$ (2)
where $y$ depends on object $x_{\texttt{obj}}$ and background $x_{\texttt{bg}}$ but is independent of noise $x_{\texttt{noise}}$ which yields $\mathbb{P}_{G_{\theta}}(y\mid x)=\mathbb{P}_{G_{\theta}}(x_{\texttt{obj}},x_{\texttt{bg}})$. $x_{\texttt{noise}}$ is the noise disturbance during satellite imaging that makes semantic confusion pan2023reducing ; pan2023prior ; pan2024pir , leading to inter-class similarity and significant intra-class variation of satellite images.
Copy-Paste Augmentation. Copy-Paste ghiasi2021simple is a data augmentation technique that involves copying objects or regions from one image and pasting them into another to create new composite scenes. As shown in Figure 3, $\texttt{Copy-Paste}(x^{a}_{obj},x^{b})$ represents the operation of copying the objects of image $x^{a}$ to image $x^{b}$. However, the Copy-Paste introduces compositional artifacts or non-smooth transitions, etc., that alter the statistical properties of the image distribution. These artifacts and transitions can typically be mitigated through the training process of DMs. More details are described in the Appendix.
4 Counterfactual Composition for Controllable Diffusion Training
Due to the inherent characteristics of satellite imagery, satellite images often exhibit high inter-class similarity and significant intra-class variation, posing challenges for RSI interpretation. To get more informative data distribution from a generative DM, we aim to approximate a real-world distribution $\mathcal{D}^{\texttt{real}}$ with as much training data distribution $\mathcal{D}^{\texttt{train}}\subset\mathcal{D}^{\texttt{real}}$ as possible. Constructing training data with diverse backgrounds and objects enables the DM to learn rich and diverse semantic information in open remote sensing scenarios. To achieve this, we enhance scene diversity through counterfactual composition, which involves combining existing object categories with diverse background contexts. The definition is as follows.
Figure 3: Left: Copy-Paste used in CF-Comp Strategy. Right: CLIP-based rule filtering retains high-quality images.
Definition 1 (Counterfactual Composition).
Given a set of source elements $A_{1},A_{2},\dots,A_{n}$}, where each $A_{i}$ represents a specific semantic component (e.g., object, region, or attribute) extracted from a distinct instance, a counterfactual sample is constructed by recombining these components as:
$x^{\prime}=\mathcal{CF}(A_{1}^{(i)},A_{2}^{(j)},\dots,A_{n}^{(k)}).\quad i\neq j\neq k$ (3)
Here, $A_{i}^{(i)}$ denotes the $i$-th component drawn from the $i$-th source instance, and $\mathcal{CF}(\cdot)$ is a composition function that logically combines elements from different instances to form a counterfactual sample $x^{\prime}$, where $x^{\prime}$ is out of the distribution of $\mathcal{D}^{\texttt{real}}$, being plausible yet not observed in reality.
Remark 1.
Counterfactual Composition allows for the generation of novel and logically consistent scenes by fusing diverse elements. The goal is to preserve semantic and structural coherence while expanding the diversity and improving the generalization of possible inputs.
Input: Training dataset $\mathcal{D}_{\texttt{train}}=\{(\mathcal{I},\mathcal{M},\mathcal{T})\}_{i=1}^{|\mathcal{D}_{\texttt{train}}|}$, where $\mathcal{I}$ is the ground-truth image, $\mathcal{M}$ is the mask prompt, and $\mathcal{T}$ is the text prompt
Output: Conditional diffusion model $\mathcal{M}_{\theta}$ with parameters $\theta$
for each training step do
   sample a mini-batch Image-Mask-Text triples $\mathcal{B}_{\texttt{ori}}=(\mathcal{I},\mathcal{M},\mathcal{T})$ from dataset $\mathcal{D}$;
 $\triangleright$ Counterfactual Composition
 $B_{\texttt{copy}}\leftarrow\{\}$, a thresholds $\alpha_{0}$, $\beta_{0}$, $\eta_{0}$
 for $(x^{a},m^{a},t^{a}),(x^{b},m^{b},t^{b})$ in $(\mathcal{I},\mathcal{M},\mathcal{T})\times(\mathcal{I},\mathcal{M},\mathcal{T})$  do
    $\alpha=\texttt{ICS}(x^{a},x^{b})$, $\beta=\texttt{MOR}(m^{a},m^{b})$, $\eta=\texttt{TSS}(t^{a},t^{b})$;
    if $\alpha$, $\beta$, $\eta$ > $\alpha_{0}$, $\beta_{0}$, $\eta_{0}$ then
       $(x^{\prime},m^{\prime},t^{\prime})=\texttt{Copy-Paste}(x^{a}_{\texttt{obj}},x^{b})$;
       $B_{\texttt{copy}}\leftarrow(x^{\prime},m^{\prime},t^{\prime})$;
  Get a new mini-batch Image-Mask-Text triples $\mathcal{B}=\mathcal{B}_{\texttt{ori}}+\mathcal{B}_{\texttt{copy}}$;
  Encode text prompt $\mathcal{T}$ using a frozen text encoder to obtain $E_{\mathcal{T}}$;
  Sample noise $\epsilon\sim\mathcal{N}(0,1)$ and timestep $t$;
  Generate noisy image: $x_{t}=\sqrt{\alpha_{t}}\cdot\mathcal{I}+\sqrt{1-\alpha_{t}}\cdot\epsilon$;
  Extract mask features $F_{\mathcal{M}}$ from $\mathcal{M}$ using $\mathcal{M}_{\theta}$ with parameters $\theta$;
  Inject mask features into a frozen UNet and predict noise $\hat{\epsilon}$:
     $\hat{\epsilon}=\text{UNet}(x_{t},t,E_{\mathcal{T}},\mathcal{M}_{\theta}(\mathcal{M}))$;
  Compute global loss and local loss: $\mathcal{L}=\|\hat{\epsilon}-\epsilon\|^{2}+\gamma\|F_{\mathcal{M}}\cdot\hat{\epsilon}-F_{\mathcal{M}}\cdot\epsilon\|^{2}$;
  Update $\theta$ using gradient descent to minimize $\mathcal{L}$;
Algorithm 1 Counterfactual Composition for Controllable Diffusion Training
Unrealistic Logical Scene. We aim to construct an unrealistic logical scene by counterfactual composition, i.e., satellite imagery of real-world data that does not exist in the real world but is logically present. This is also intended to maintain consistency in the image-label union space and prevent disruptions caused by arbitrary counterfactual composition. As shown in Figure 2, we use the Copy-Paste to perform counterfactual composition. For example, the unrealistic logical scene is obtained by combining the two images’ small and large vehicle objects in Figure 3. An unrealistic logical scene can be judged from three dimensions: channel, pixel, and semantic space. We define three criteria to measure whether two images can be combined into an unrealistic logical scene:
From the channel space dimension, the Image Color Sensitivity (ICS) can be obtained as
$\texttt{ICS}(x_{a},x_{b})=\begin{cases}\mathbf{1}_{\left\{\left|S(x^{a})-S(x^{b})\right|<s_{0}\right\}},&\text{if }C_{a}=C_{b}\\ 0,&\text{if }C_{a}\neq C_{b}\end{cases},$ (4)
where the color sensitivity $S=\operatorname{Var}(R-G)+\operatorname{Var}(R-B)+\operatorname{Var}(G-B)$, $C_{a}$ and $C_{b}$ are the number of channels of satellite images $x_{a}$ and $x_{b}$ respectively.
From the pixel space dimension, the Mask Overlap Rate (MOR) can be obtained as
$\texttt{MOR}(m^{a},m^{b})=\begin{cases}\dfrac{|m^{a}\cap m^{b}|}{|m^{a}\cup m^{b}|},&\text{if }|m^{a}\cup m^{b}|>0\\ 0,&\text{otherwise}\end{cases}.$ (5)
From the semantic space dimension, the text semantic similarity (TSS) can be obtained as
$\texttt{TSS}(t^{a},t^{b})=\frac{t^{a}\cdot{t^{b}}^{\top}}{\|t^{a}\|\cdot\|t^{b}\|}.$ (6)
Input: Pre-trained conditional diffusion model $\mathcal{M}_{\theta}$ parameters $\theta$, sampling steps $T$, sampling condition set $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$
Output: Synthetic dataset $\mathcal{S}=\{x_{0}^{(i)},c_{i}\}_{i=1}^{|\mathcal{C}|}$
Initialize empty synthetic dataset: $\mathcal{S}\leftarrow\{\}$;
for each condition $c_{i}\in\mathcal{C}$ do
   Sample initial noise $x_{T}^{(i)}\sim\mathcal{N}(0,\mathbf{I})$;
  Encode text prompt $\mathcal{T}$ using a frozen text encoder to obtain $E_{\mathcal{T}}$;
  Extract mask features $F_{\mathcal{M}}$ from $\mathcal{M}$ using $\mathcal{M}_{\theta}$ with parameters $\theta$;
 for $t=T$ down to 11 do
      Predict noise: $\hat{\epsilon}_{t}=\text{UNet}(x_{t}^{(i)},t,E_{\mathcal{T}},\mathcal{M}_{\theta}(c_{i}))$;
     Update latent variable using reverse diffusion equation:
        $x_{t-1}^{(i)}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}^{(i)}-\sqrt{1-\alpha_{t}}\cdot\hat{\epsilon}_{t}\right)+\sigma_{t}z$,
        where $z\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $z=0$;
 $\triangleright$ Rule-based Filtering
 if $Score_{\texttt{CLIP}}(x)>S_{0}\mid Score_{\texttt{CLIP}}(x_{obj})>S_{0}$ then
     Add generated sample to dataset: $\mathcal{S}\leftarrow\mathcal{S}\cup\{(x_{0}^{(i)},c_{i})\}$;
return Final dataset $\mathcal{S}$ sampled from conditional distribution $\mathcal{X}_{\mathcal{C}}$
Algorithm 2 Training Data Synthesis
Mixed Data Distribution. In real remote sensing scenarios, the object and background of an image typically follow a consistent data distribution. However, in unrealistic logical data distributions, discrepancies between foreground and background distributions often lead to out-of-distribution hsu2020generalized . By integrating real-world and unrealistic logical data distributions, the latent data manifold becomes more complex and diverse, allowing the model to generalize better to unseen scenarios while preserving the essential characteristics of the original distribution. Assuming that the object and background of the image can be modeled as Gaussian distributions, respectively, as:
$x_{\texttt{obj}}\sim\mathcal{N}(\mu_{\texttt{obj}},\sigma^{2}_{\texttt{obj}}I),\quad x_{\texttt{bg}}\sim\mathcal{N}(\mu_{\texttt{bg}},\sigma^{2}_{\texttt{bg}}I),$ (7)
where $(\mu_{\texttt{obj}},\sigma^{2}_{\texttt{obj}})$ and $(\mu_{\texttt{bg}},\sigma^{2}_{\texttt{bg}})$ denote the respective mean and variance of the object and background distributions. Assuming independence and linearity of expectation, the mean and variance of the resulting image $x^{\prime}$ with mask $m$ by Copy-Paste can be derived as:
$\mu_{\texttt{mix}}$ $=\alpha\mu_{\texttt{obj}}+(1-\alpha)\mu_{\texttt{bg}}$ (8)
$\sigma^{2}_{\texttt{mix}}$ $=\alpha\sigma^{2}_{\texttt{obj}}+(1-\alpha)\sigma^{2}_{\texttt{bg}}+\alpha(1-\alpha)(\mu_{\texttt{obj}}-\mu_{\texttt{bg}})^{2},$
where $\alpha=\frac{1}{HW}\sum_{i,j}m_{i,j}$ represents the proportion of foreground pixels in the image. The additional term $\alpha(1-\alpha)(\mu_{\texttt{obj}}-\mu_{\texttt{bg}})^{2}$ in the variance reflects the distributional shift caused by the mismatch between object and background statistics. CF-Comp allows models to learn richer and more informative representations by controlling the Copy-Paste ratio $\Sigma\alpha$ in remote sensing. We note that in satellite images for background distraction and intra-class variability ma2024direction , $\mu_{\texttt{obj}}\approx\mu_{\texttt{bg}}$ and $\sigma^{2}_{\texttt{obj}}\approx\sigma^{2}_{\texttt{bg}}$, but in natural scene images, these parameters may differ significantly, resulting in distributional shifts huang2021importance .
Global and Local Loss. Unlike binary masks, semantic masks embed category-specific information, making pixel-level precision essential. However, using a global constraint alone is inadequate for modeling such fine-grained control. Therefore, we integrate global and local constraints to achieve a more accurate generation. We extract semantic mask features $F_{\mathcal{M}}$ from mask $\mathcal{M}$ using condition model $\mathcal{M}_{\theta}$ with parameters $\theta$ and inject semantic mask features into a frozen UNet ronneberger2015u and predict noise $\hat{\epsilon}$:
$\mathcal{L_{\texttt{global}}}=\|\hat{\epsilon}-\epsilon\|^{2},\quad\mathcal{L_{\texttt{local}}}=\|F_{\mathcal{M}}\cdot\hat{\epsilon}-F_{\mathcal{M}}\cdot\epsilon\|^{2},\quad\mathcal{L}=\mathcal{L_{\texttt{global}}}+\gamma\mathcal{L_{\texttt{local}}}$ (9)
where $\gamma$ is the local constraint factor.
Training Process. Algorithm 1 shows the batch-based CF-Comp method for condition diffusion EarthSynth training. We follow Zhang et al. zhang2023adding to train a conditional diffusion model.
Method Scene Classification
RSICD* (Top1 / Top5) DIOR (Top1 / Top5)
Txt2Img-MHN(VQVAE)† xu2023txt2img 32.7 / 75.5 - / -
Txt2Img-MHN(VQGAN)† xu2023txt2img 40.9 / 72.7 - / -
CRS-Diff† tang2024crs 57.1 / 79.0 - / -
StableDiffusion† rombach2022high 61.3 / 88.3 41.5 / 73.0
InstanceDiffusion† wang2024instancediffusion 59.1 / 88.1 44.5 / 79.0
ControlNet† zhang2023adding 55.5 / 85.5 46.5 / 78.5
EarthSynth (Ours) 60.0 / 91.8 49.0 / 80.0
Table 1: CLIP-based scene classification accuracy on RSICD and DIOR datasets with Acc. †: training on remote sensing data.
Method Data Usage Object Detection
DOTAv2 DIOR
Base GroundingDINO Real 56.3 74.0
+ StableDiffusion rombach2022high Real + Synth - -
+ ControlNet zhang2023adding Real + Synth 57.4 (+1.1) 74.1 (+0.1)
+ EarthSynth (Ours) Real + Synth 58.4 (+2.1) 74.3 (+0.3)
Table 2: Object detection on DOTAv2 and DIOR dataset with mAP, validated on the open-vocabulary object detection task.
5 Training Data Synthesis
Since the quality of data generated by DMs can vary significantly, we propose a rule-based filtering method, R-Filter, to further refine the generated samples and retain only those that meet predefined quality criteria. Algorithm 2 shows training data synthesis with EarthSynth.
Condition Sampling. During the data synthesis stage, conditions $c_{i}=(m_{i},t_{i})$ are randomly sampled by category from the training condition set $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$. We also apply random rotation, scaling, and merging, based on the mask and text prompts, to get more diverse conditions. Figure 5 shows that EarthSynth can generate some unrealistic logical scenes controlled by different text prompts.
Label Mapping. Different label mapping functions $\mathcal{F}$ are employed for different downstream tasks. For scene classification, the category label $y_{c}$ is directly obtained from the associated text $t_{i}$. For semantic segmentation, the semantic mask $m_{i}$ and the category label $y_{c}$ are derived by extracting class mappings from the corresponding mask. For object detection, bounding boxes are generated by extracting the contour features of the masks by using the Ramer-Douglas-Peucker cao2022automatic algorithm.
Rule-based Filtering. We propose R-Filter, a rule-based method that uses CLIP scores to evaluate {image $x$, object $x_{obj}$, background $x_{bg}$} triplets by computing overall, object-specific, and background scores, as shown in Figure3. Since both $x_{obj}$ and $x_{bg}$ are related to the label $y$, we retain samples with high overall or object-specific scores for training downtown models by setting the CLIP score threshold $S_{0}$.
6 Experiment
In this section, we evaluate EarthSynth on scene classification, object detection, and semantic segmentation, including performance analysis, ablation studies, and visualization analysis.
EarthSynth-180K. The diffusion model is trained on cross-satellite and multi-sensor data to enhance generation diversity and improve object modeling across varying observation conditions. EarthSynth-180K is built from OEM xia2023openearthmap , LoveDA wang2021loveda , DeepGlobe demir2018deepglobe , SAMRS wang2023samrs , and LAE-1M pan2025locate datasets from different satellites and sensors, and enhanced with mask and text prompts. We leverage many pixel-level mask annotations and semantic-level texts as prompts for DMs. By applying Random Cropping Strategy and Category-Augmentation Strategy to the EarthSynth-180K dataset, we obtain about 180K high-quality triplets consisting of images, semantic masks, and texts. This dataset has a wide range of categories and is labeled with semantic segmentation, which can be used to improve object detail reconstruction and category understanding in DMs. More details can be found in the Appendix.
Evaluation. We use multiple tasks to evaluate data generation capability, including scene classification cheng2017remote , object detection zou2023object , and semantic segmentation guo2018review . We construct an evaluation task that progresses from coarse-grained to fine-grained levels, spanning from image-level to pixel-level, to assess the generalization capability of synthetic data.
Experiment Setup. All experiments are performed using four NVIDIA A100 GPUs, and the complete training of EarthSynth requires approximately 4 $\times$ 45 GPU hours. EarthSynth is initialized with the pretrained Stable Diffusion v1-5 rombach2022high weights. Training uses mixed precision to improve computational efficiency and reduce memory consumption. A batch size of 8 per device is used, with gradient accumulation over four steps, resulting in an adequate batch size of 32. In the CF-Comp setting, we set $s_{0}=150$, $\alpha_{0}=1$, $\beta_{0}=0.02$, and $\eta_{0}=0.6$. The local constraint $\gamma$ is set to 10. The training runs for 40,000 steps. A constant learning rate of 1e-5 is adopted without any warm-up phase, and gradient clipping with a maximum norm of 1 is applied to ensure training stability. To improve data quality, we use the CLIP-ViT-B/32 model radford2021learning with the CLIP score threshold $S_{0}$ set to 0.4.
6.1 Comparative Results
Method Data Usage mAP
1-shot Detic zhou2022detecting Real 4.1
DE-ViT zhang2023detect Real 14.7
CD-ViTO fu2024cross Real 17.8
GroundingDINO liu2024grounding Real 11.7
ETS pan2025enhancesearchaugmentationsearchstrategy Real 12.7
+ ControlNet zhang2023adding Synth 9.2
+ ControlNet zhang2023adding Real + Synth 13.1
+ EarthSynth (Ours) Synth 9.5
+ EarthSynth (Ours) Real + Synth 13.9
5-shot Detic zhou2022detecting Real 12.1
DE-ViT zhang2023detect Real 23.4
CD-ViTO fu2024cross Real 26.9
GroundingDINO liu2024grounding Real 27.7
ETS pan2025enhancesearchaugmentationsearchstrategy Real 29.3
+ ControlNet zhang2023adding Synth 23.0
+ ControlNet zhang2023adding Real + Synth 33.9
+ EarthSynth (Ours) Synth 23.6
+ EarthSynth (Ours) Real + Synth 34.1
10-shot Detic zhou2022detecting Real 15.4
DE-ViT zhang2023detect Real 25.6
CD-ViTO fu2024cross Real 30.8
GroundingDINO liu2024grounding Real 36.4
ETS pan2025enhancesearchaugmentationsearchstrategy Real 37.5
+ ControlNet zhang2023adding Synth 25.3
+ ControlNet zhang2023adding Real + Synth 40.2
+ EarthSynth (Ours) Synth 26.4
+ EarthSynth (Ours) Real + Synth 40.7
Table 3: The few-shot detection results on the DIOR dataset.
We evaluate downstream remote sensing tasks in open-vocabulary scenarios, including scene classification, object detection, and semantic segmentation. Training data is synthesized using remote sensing-specific methods trained on RSICD lu2017exploring Txt2Img-MHN xu2023txt2img , CRS-Diff tang2024crs , and baseline methods trained on EarthSynth-180K, Stable Diffusion rombach2022high , InstanceDiffusion wang2024instancediffusion , ControlNet zhang2023adding . We adopt open-vocabulary methods, CLIP radford2021learning , GroundingDINO liu2024grounding , and GSNet ye2025towards for downtown evaluation method. Real denotes real-world data, and Synth refers to diffusion-generated data. Note that most existing remote sensing diffusion models are based on the above architectures, lack category control, and are not optimized for open-vocabulary understanding tasks. More detailed analyses are provided in the Appendix.
Scene Classification. We uniformly generated 10 images per category and averaged the results over three runs for evaluation. Table1 presents the comparative results of CLIP-based scene classification regarding Top-1 and Top-5 accuracy on RSICD lu2017exploring and DIOR li2020object datasets. RSICD* is a subset of the original RSICD dataset containing 11 classes, used to align different dataset settings of different methods that do not include the RSICD dataset. Compared to VAE-based and GAN-based methods, diffusion-based approaches exhibit a clear advantage in generation quality. On RSICD, our method ranks second in Top-1 accuracy and first in Top-5 accuracy for classification. It also outperforms the baseline ControlNet by 6.3. Our method achieves superior scene classification performance on DIOR, indicating its ability to generate images with higher classification scores. These results suggest the potential of our approach in downstream tasks.
Object Detection. We evaluate the effectiveness of our proposed method on object detection by training the GroundingDINO model on DOTAv2 or DIOR dataset with 256 synthetic images per category. Table 2 experiments on the DOTAv2 xia2018dota and DIOR li2020object datasets show that our approach outperforms ControlNet, achieving improvements of 2.1 and 0.3 compared to training solely on real data. We also experimented with a few-shot object detection. Table 3 shows that ETS pan2025enhancesearchaugmentationsearchstrategy consistently outperforms close-source baselines across all shot settings. ETS achieves 37.5 mAP in the 10-shot setting, surpassing Detic at 15.4 and DE-ViT at 25.6. Introducing synthetic data leads to notable improvements, especially when combined with real data. In the 5-shot setting, ETS improves from 29.3 using only real data to 33.9 with ControlNet-generated data and 34.1 with EarthSynth-generated data. These results show that adding synthetic data from EarthSynth can improve open-vocabulary object detection.
Semantic Segmentation. We generated 256 synthetic images for each category and combined them with the real EarthSynth-180K dataset to train the open-vocabulary segmentation method GSNet. Table 4 reports the mIoU results on four unseen semantic segmentation datasets: Potsdam ISPRS2013 , FloodNet rahnemoonfar2021floodnet , FAST wang2023samrs , and FLAIR garioud2023flair . The baseline GSNet model achieves solid performance using only real data, particularly on the Potsdam dataset, and incorporating synthetic data from ControlNet results in mixed outcomes, with performance improvements on FloodNet by 5.4, FAST by 0.9, and FLAIR by 1.1. Still, a degradation of Potsdam by 5.2, suggesting potential issues related to multi-source data alignment. In contrast, EarthSynth consistently improves performance across all datasets, with gains of 2.1 on Potsdam, 10.5 on FloodNet, 0.6 on FAST, and 2.3 on FLAIR. These results demonstrate the effectiveness of EarthSynth-generated data in enhancing semantic segmentation under diverse scenarios.
Method Data Usage Semantic Segmentation
Potsdam FloodNet FAST FLAIR
Base GSNet Real 40.6 33.9 16.8 19.3
+ StableDiffusion rombach2022high Real + Synth - - - -
+ ControlNet zhang2023adding Real + Synth 35.4 (-5.2) 39.3 (+5.4) 17.7 (+0.9) 20.4 (+1.1)
+ EarthSynth (Ours) Real + Synth 42.7 (+2.1) 44.4 (+10.5) 17.4 (+0.6) 21.6 (+2.3)
Table 4: Semantic segmentation on Potsdam, FloodNet, FAST, and FLAIR datasets with mIoU.
Figure 4: Effect of samples per class on DOTAv2 dataset.
Method mAP
*only real 56.3
baseline -
+ $\mathcal{L}_{\text{local}}$ 56.5
+ $\mathcal{L}_{\text{local}}$ + R-Filter 57.4
+ CF-Comp -
+ CF-Comp + $\mathcal{L}_{\text{local}}$ 57.9
+ CF-Comp + $\mathcal{L}_{\text{local}}$ + R-Filter 58.4
Table 5: Detection performance on DOTAv2 using different modules configurations. “–” is a failure to learn mask-based semantic control for labeled sample generation.
6.2 Ablation Studies
In this section, we perform the ablation studies of the submodule, sample size, and CF-Comp.
Submodule. We fix the random seed during inference to ensure consistency across comparative experiments, guaranteeing that identical templates are used throughout the evaluation. As shown in Table5, object detection performance on DOTAv2 using different submodule configurations. We found that introducing the local loss $\mathcal{L}_{\texttt{local}}$ can learn layout control of the semantic mask. Adding R-Filter further boosts performance to 57.4. Incorporating CF-Comp alongside $\mathcal{L}_{\texttt{local}}$ brings a larger gain, reaching 57.9. The full configuration with CF-Comp, $\mathcal{L}_{\text{local}}$, and R-Filter achieves the best result of 58.4 mAP. CF-Comp and R-Filter effectively retain higher-quality informative samples for downstream tasks.
Sample Size. Figure 4 shows that performance improves as the number of synthetic samples per class increases, reaching the highest mAP of 58.4 at 128 samples per class. However, further increases lead to marginal declines, suggesting that moderate-scale synthesis is most effective, while excessive data may introduce redundancy or noise.
CF-Comp and CLIP of R-Filter. Further ablation studies on CF-Comp and CLIP of R-Filter are in the Appendix.
6.3 Visualization Analysis
Synthesis. Figure 5 shows the synthesis satellite images with FID scores heusel2017gans on DOTAv2 dataset. Compared to the baseline ControlNet, EarthSynth in larger distributional shifts and generates higher-quality images. In addition, by combining prompts with transformed masks, we can produce more robust images for downstream tasks.
Embedding. Figure 6 shows the t-SNE van2008visualizing distributions and FID scores for three representative categories, illustrating the alignment between synthetic and real data. The use of CF-Comp increases sample diversity but also leads to larger distributional shifts, as indicated by higher FID scores. The introduction of R-Filter also increases this gap, with FID scores consistently increasing across three categories. These results suggest that CF-Comp and R-Filter enhance the semantic diversity and distributional spread of synthetic samples, potentially improving generalization for downstream tasks.
Figure 5: Left: Visualization of synthesis satellite images on DOTAv2 dataset. Right: EarthSynth can generate some unrealistic logical scenes controlled by different text prompts.
Figure 6: Embedding visualizations for some categories on the EarthSynth-180K dataset.
7 Conclusion
We propose EarthSynth, a diffusion-based generative foundation model trained on the EarthSynth-180K dataset, which is the first large-scale remote sensing dataset for diffusion training. This work addresses the challenge of poor generalization in task-oriented synthesis for RSI interpretation tasks. To balance layout controllability and category diversity during training, EarthSynth adopts the batch-based CF-Comp strategy, enabling precise layout control for RSI generation. Additionally, it incorporates the R-Filter post-processing method to extract more informative synthesized samples for downstream tasks. EarthSynth is evaluated in open-world scenarios, offering a practical and scalable solution to advance RSI interpretation through synthetic data.
Limitations. Our work focuses on RGB-based remote sensing imagery, excluding standard multispectral data. It also incurs higher training costs, a known challenge in image generation. Further discussion is provided in the Appendix.
Appendix A Technical Appendices and Supplementary Material
A.1 More Preliminaries
A.1.1 Copy-Paste Augmentation.
Copy-Paste ghiasi2021simple is a data augmentation technique that involves copying objects or regions from one image and pasting them into another to create new composite scenes. $\texttt{Copy-Paste}(x^{a}_{obj},x^{b})$ represents the operation of copying the objects of image $x^{a}$ to image $x^{b}$. However, the Copy-Paste introduces compositional artifacts or non-smooth transitions, etc., that alter the statistical properties of the image distribution. These artifacts and transitions can typically be mitigated through the training process of DMs. Given two satellite images, $x^{a}$ and $x^{b}$ with their masks $m^{a}$ and $m^{b}$ and text embddings $t^{a}$ and $t^{b}$, the $\texttt{Copy-Paste}(x^{a},x^{b})$ operation can be defined as follows:
$x^{\prime}=x^{a}+\mathbf{1}_{\{m^{a}=0\}}\cdot x^{b},$
$m^{\prime}=m^{a}+\mathbf{1}_{\{m^{a}=0\}}\cdot m^{b},$
$t^{\prime}=t^{b}+t^{a}\setminus(t^{a}\cap t^{b}),$
where $\mathbf{1}_{\{m^{a}=0\}}$ denotes an indicator function that returns 1 when $m^{a}=0$, and 0 otherwise. $\texttt{Copy-Paste}(x^{a}_{obj},x^{b})$ represents the operation of copying the objects of image $x^{a}$ to image $x^{b}$.
A.2 EarthSynth-180K Dataset
EarthSynth-180K is derived from OEM, LoveDA, DeepGlobe, SAMRS, and LAE-1M datasets from different satellites. The satellite sources of the EarthSynth-180K dataset are shown in Table 7. It is further enhanced with mask and text prompt conditions, making it suitable for training foundation DMs. The EarthSynth-180K dataset is constructed using the Random Cropping and Category-Augmentation strategies. The category distribution of the EarthSynth-180K dataset is presented in Figure 7, along with the corresponding category-to-abbreviation mapping shown in Table 6. Although remote sensing focuses on a limited number of classes, this study validates the feasibility of the method on major classes by expanding the vocabulary, and the limited number of classes does not affect the conclusions.
Figure 7: Category distribution of the EarthSynth-180K.
A.2.1 Random Cropping Strategy
To standardize the input resolution for the DM, we employ a random cropping strategy to generate $\times512$ image patches. For lengths smaller than 1024, resample to 512; the insufficient edge parts are filled with zero values. The same cropping operation is applied to the corresponding semantic masks to maintain spatial consistency. Specifically, if the image size is larger than the target crop size, a random top-left coordinate $(x,y)$ is selected to extract a patch of the desired dimensions. If the image is smaller than the crop size, no cropping is performed, and the region starting from the top-left corner is used directly. To construct textual descriptions for the categories, we employ a template-based approach by appending category names to the phrase “A satellite image of [class1], [class2], …”, resulting in descriptions such as “A satellite image of swimming pool, ship, small vehicle, harbor”. We construct around 180K triplets of images, textual descriptions, and semantic masks. Note that center-based cropping is not required, as the categories in the dataset are heterogeneous and spatially scattered.
Figure 8: A category-augmentation strategy to construct multiple object-background pairs.
A.2.2 Category-Augmentation Strategy
We apply data augmentation to each image to enhance the model’s understanding of individual categories and enable fine-grained, category-specific control during generation, creating multiple distinguishable foreground-background pairs. Additionally, this approach helps improve the probability of sample combinations in the batch-based CF-Comp strategy. Figure 8 shows that these augmented triplets are generated from the previously obtained image-text-mask triplets by isolating non-background categories in the original masks and corresponding textual descriptions, resulting in a set of new, category-focused triplets. We obtained approximately 500K paired textual descriptions and semantic masks of varying granularities through category augmentation.
Category Abbr. Category Abbr. Category Abbr.
background BG ground track field GF vehicle VE
bare land BL small vehicle SV windmill WM
grass GR baseball diamond BD expressway service area EA
pavement PA tennis court TC expresswalltoll station ET
road RO roundabout RA dam DM
tree TR storage tank ST golf field GO
water WA harbor HA overpass OP
agriculture land AL container crane $C$ stadium SD
buildings BU airport AP train station TS
forest land FL helipad HP large vehicle LV
barren land BA chimney CH swimming pool SP
urban land UL helicopter HE bridge BR
plane PL ship SH soccer ball field SF
basketball court BC
Table 6: The main category to abbreviation mapping.
Dataset Data Sources Sensor Type
OEM xia2023openearthmap Existing Benchmark Dataset, Various Satellite Operators and Agencies Satellite, Aircraft, and UAV
LoveDA wang2021loveda Google Earth Platform Satellite
DeepGlobe demir2018deepglobe WorldView-2 Satellite
SAMRS wang2023samrs Sentinel-1, Sentinel-2, PlanetScope, and others Satellite, Aircraft, and UAV
LAE-1M pan2025locate Existing Object Detection Datasets, Google Earth Platform Satellite, Aircraft, and UAV
Table 7: Data sources and sensor type of EarthSynth-180K.
A.2.3 Data Sources of EarthSynth-180K
Table 7 summarizes the satellite platforms and corresponding sensor types that contribute to the EarthSynth-180K dataset.
OEM. The Open Earth Map (OEM) dataset is a global initiative to advance open machine learning-based mapping techniques using remote sensing data. It focuses on extracting semantic and height information, such as land cover maps and digital elevation models (DEMs), to support environmental monitoring, urban planning, and disaster management applications. The dataset comprises 5,000 aerial and satellite images with manually annotated 8-class land cover labels at a 0.25–0.5m ground sampling distance, covering 97 regions from 44 countries across six continents.
LoveDA. LoveDA dataset is designed for land-cover domain adaptation semantic segmentation. It contains 5,987 high spatial resolution (0.3m) remote sensing images from three cities in China, including urban and rural scenes. The images are sourced from the Google Earth Platform, providing real-world urban and rural remote sensing images for semantic segmentation and unsupervised domain adaptation tasks.
DeepGlobe. DeepGlobe dataset is part of the DeepGlobe 2018 Satellite Image Understanding Challenge, which includes three public competitions for segmentation, detection, and classification tasks on satellite images. The dataset consists of high-resolution satellite images with a 50cm pixel resolution collected by DigitalGlobe’s WorldView series satellites. It is used for road extraction and building detection tasks.
Figure 9: Some images of the EarthSynth-180K dataset are derived from SAR imagery.
SAMRS. SAMRS dataset is a large-scale remote sensing segmentation dataset developed using the Segment Anything Model kirillov2023segment . It leverages existing remote sensing object detection datasets to generate a comprehensive dataset for semantic segmentation, instance segmentation, and object detection tasks. The dataset comprises 105,090 images with 1,668,241 instances, surpassing existing high-resolution remote sensing segmentation datasets in size by several orders of magnitude. It integrates data from various sources, including Sentinel-1, Sentinel-2, and PlanetScope satellites.
LAE-1M. LAE-1M dataset is a remote sensing object detection dataset with broad category coverage. It is constructed by unifying up to 10 remote sensing datasets to create a comprehensive collection for open-vocabulary object detection tasks. The dataset includes high-resolution optical satellite imagery, though specific satellite sources are not disclosed. We used part of the LAE-1M dataset as expanded semantic diversity. While this work mainly concentrates on optical image understanding and generation, a portion of the ship category in the dataset is derived from SAR imagery, as shown in Figure 9. Due to the scarcity of such data, understanding cross-sensor categories remains a vital aspect to consider.
Method Condition Type
Semantic-level (class, text) Region-level (box) Pixel-level (mask, sketch)
Text2Earth liu2025text2earth
DiffusionSat khanna2023diffusionsat
GeoSynth sastry2024geosynth
CRS-Diff tang2024crs
SatSynth toker2024satsynth
AeroGen tang2024aerogen
MMO-IG yang2025mmo
EarthSynth (Ours)
Table 8: Comparison of remote sensing diffusion methods based on different levels of prompts.
Type Method Training Dataset FID Score CLIP Score
Single
StableDiffusion rombach2022high RSICD 103.4 26.1
StableDiffusion rombach2022high DIOR 228.1 26.2
InstanceDiffusion wang2024instancediffusion RSICD 138.6 24.7
Text2Earth liu2025text2earth Git-10M 24.5* -
DiffusionSat khanna2023diffusionsat fMoW 15.8* 17.2*
GeoSynth sastry2024geosynth US Cities 12.3* / 171.1 30.3* / 25.0
SatSynth toker2024satsynth iSAID, LoveDA, OEM - -
MMO-IG yang2025mmo DIOR 34.5* -
AeroGen tang2024aerogen HRSC,DIOR 38.6* -
Multi
CRS-Diff tang2024crs RSICD, fMoW 50.7* / 107.0 20.3* / 23.6
InstanceDiffusion wang2024instancediffusion RSICD, RSITMD, UCMerced 123.1 25.0
ControlNet zhang2023adding EarthSynth-180K 183.8 25.9
EarthSynth (Ours) EarthSynth-180K 198.7 26.1
Table 9: Comparison of FID and CLIP scores trained on single-source (Single) and multi-source Multi data. We calculated the FID score between the generated images and the RSICD dataset. * is from the original papers.
A.3 Remote Sensing Diffusion Models
Table 8 compares remote sensing diffusion models across three prompt levels: semantic, region, and pixel. Most methods support only semantic-level prompts, offering global but coarse control. Some models introduce region-level prompts using bounding boxes to enhance spatial precision. Pixel-level prompts, such as masks and sketches, provide the most detailed control and are used by CRS-Diff, GeoSynth, and EarthSynth. EarthSynth uniquely combines semantic and pixel-level prompts, enabling high-level semantics and fine-grained spatial guidance. This reflects a shift toward more precise and controllable image generation.
High-Resolution Synthesis Text2Earth, DiffusionSat, and GeoSynth are representative methods to generate high-resolution satellite imagery. These models leverage diffusion-based generative frameworks or text-to-image architectures to reconstruct fine-grained spatial details, often guided by auxiliary inputs such as text descriptions, semantic maps, or multi-modal signals. Their primary applications lie in image restoration, super-resolution, cloud removal, and spectral enhancement, which are essential for improving satellite data’s visual and analytical quality in scientific and operational settings.
Method Visual Quality Semantic Richness Overall Score
ControlNet 79.51 23.76 51.64
EarthSynth (Ours)) 80.19 25.95 53.07
Table 10: Image scoring results on 100 diffusion-generated images using GPT-4.
Task-Oriented Synthesis CRS-Diff, SatSynth, AeroGen, and MMO-IG are designed with downstream utility, focusing on generating synthetic data tailored for specific tasks such as land cover classification, object detection, and change detection. These models incorporate task-specific priors, including class labels, semantic layouts, or instance-level masks, to guide the generation process. By aligning the synthesized data with the needs of target tasks, these methods enhance model generalization in low-resource scenarios, enable domain adaptation, and facilitate pretraining or fine-tuning of models in remote sensing applications.
A.4 More Experiments
A.4.1 Comparison of FID and CLIP scores
Table 9 compares fake data from different models using FID and CLIP scores. FID measures distribution distance to all images of the RSICD dataset, while CLIP evaluates semantic alignment. The results reveal a mismatch between FID and CLIP metrics across models. For example, GeoSynth shows low FID (12.3) and high CLIP (30.3), indicating strong visual and semantic quality. In contrast, StableDiffusion on RSICD or DIOR has a high FID (103.4, 228.1) but is similar to CLIP on EarthSynth (26.1). InstanceDiffusion improves FID with multi-source data, yet its CLIP score stays nearly the same (25.0 vs. 24.7), underscoring a gap between visual fidelity and semantic alignment. This suggests that FID alone is not reliable for assessing task-specific data quality.
A.4.2 Multi-modal LLM for Image Quality Evaluation
Image Scoring based on Multi-modal LLM. We use GPT-4 as a text-based tool to evaluate the generated images, since it can’t analyze pixels directly. We extract basic data like resolution, color mode, and estimated visual complexity based on color count, then turn this into a descriptive prompt that summarizes the image’s main features. Our prompt consists of two components: (1) a system prompt that directs GPT-4 as an expert evaluator capable of judging image quality and semantic richness based only on textual input. The system prompt states: “You are an expert image evaluator who assesses image quality and semantic richness from detailed descriptions alone. You do not require access to the actual image; provide a numeric score from 0 to 100 based solely on the description.” (2) a user prompt that supplies the image description and explicitly instructs GPT-4 to respond in the exact format: “Score: X. Reason: <brief explanation>”, where X is an integer between 0 and 100. The prompt explicitly forbids disclaimers such as “I am an AI and cannot view images.”
You are an expert image evaluator. Based ONLY on the description below,
rate the image on two aspects:
1. Visual Quality (integer score 0-100)
2. Semantic Richness (integer score 0-100)
Image description: <image description text>
Respond ONLY with exactly the following format and nothing else:
Visual Quality Score: <integer 0-100>
Semantic Richness Score: <integer 0-100>
Reason: <one brief sentence>
Image scoring is performed along two key dimensions: (1) visual quality, encompassing factors such as image clarity, resolution, and compositional coherence; and (2) semantic richness, including the quantity of meaningful elements, scene complexity, and the depth of emotional or narrative content. This structured prompt formulation promotes consistent and parsable outputs, reducing subjective variance. Numeric scores are extracted via regular expression matching, and average scores are computed over the entire image set to yield final metrics. The prompts are listed below:
Method CLIP Train Data Top1 Acc
Baseline (EarthSynth) - 49.07
+ RemoteCLIP RET-3 + DET-10 + SEG-4 48.13
+ CLIP Web 50.47
Table 11: CLIP-based scene classification performance on DOTA-v2 using different module configurations.
Experimental Setup and Results. We evaluated 100 generated images for each method. The assessment was performed using GPT-4, which rated the images along two dimensions: Visual Quality, reflecting the perceptual fidelity and aesthetic appeal, and Semantic Richness, indicating the depth and variety of semantic content present in the image. The Overall Average Score was computed as the arithmetic mean of the two individual scores.
As shown in Table 10, EarthSynth achieved a slightly higher visual quality score (80.19) than ControlNet (79.51), indicating marginally better perceptual quality. EarthSynth produced images with greater semantic richness, scoring 25.95 compared to 23.76 from ControlNet. Consequently, ControlNet achieved a higher overall average score of 53.07, compared to 51.64 for EarthSynth. These results suggest EarthSynth performs better at both visual quality and semantic richness. We believe its scoring results can only serve as a reference and are intended to provide one of the methods for diversified image quality evaluation.
Figure 10: EarthSynth over time-step training process.
A.4.3 Ablation Studies
Local Loss. We qualitatively analyze the role of local constraints by visualizing the training process of the EarthSynth model. As shown in Figure 11 and Figure 10, the loss function is defined as,
$\mathcal{L}=\mathcal{L}_{\texttt{global}}+\gamma\mathcal{L}_{\texttt{local}}.$
Setting $\gamma=10$ accelerates convergence and helps the model capture semantic mask information better. Compared to not using $\mathcal{L}_{\texttt{local}}$, incorporating it enables more effective layout control. We found it challenging to learn layout control without local constraints on the mask. For comparison fairness, ControlNet with $\mathcal{L}_{\texttt{local}}$ is used in object detection and semantic segmentation.
Figure 11: EarthSynth over time-step training process.
CF-Comp. As illustrated in Figure 11 and Figure 10, we qualitatively analyze the impact of local constraints by visualizing the training process of the EarthSynth model. A visual ablation study on ICS and TSS further highlights their complementary roles within our framework. Specifically, incorporating ICS encourages the model to generate more diverse and novel compositional combinations, enhancing creativity and variety. In contrast, applying TSS fosters the generation of semantically coherent and realistic compositions, improving overall plausibility. These findings also validate the effectiveness of our proposed EarthSynth with CF-Comp strategy, demonstrating its ability to balance novelty with semantic fidelity.
CLIP of R-Filter. Table 11 presents the Top-1 accuracy of different configurations of the EarthSynth model on the DOTA-v2 scene classification task. The baseline EarthSynth model achieves an accuracy of 49.07. When integrated with RemoteCLIP, performance slightly drops to 48.13, suggesting that RemoteCLIP may not effectively enhance EarthSynth in this context, possibly due to misalignment with remote sensing imagery. In contrast, incorporating the standard CLIP module leads to a significant improvement, reaching the highest accuracy of 50.47. This result highlights CLIP’s strong capability to boost scene classification performance. Also, this reacts to the semantic bias that RemoteCLIP carries in the remote sensing domain, which is not well used for filtering.
Figure 12: Examples of generated segmentation data.
A.5 More Visualization
A.5.1 Examples of EarthSynth-generated Data
Figure 12 and Figure 13 illustrate representative examples of the synthetic semantic segmentation and object detection datasets generated using the proposed EarthSynth framework. These visualizations show that the semantic segmentation labels exhibit higher precision and consistency across diverse land cover types. This can be attributed to the pixel-level supervision involved in the segmentation process, which enables more accurate delineation of object boundaries. Object detection annotations are less reliable because they rely on post-processing techniques like edge detection and bounding box generation. These methods often introduce noise or result in misalignment with the actual object extents, thereby reducing the overall annotation quality. This contrast highlights the relative robustness of the segmentation outputs and suggests that EarthSynth is particularly well-suited for applications where spatial accuracy and detailed contextual information are essential.
A.5.2 Comparison of Generated Images from Remote Sensing Diffusion Models
Figure 14 compares generated images from remote sensing diffusion models. The focus of the different methods for generating images is described below.
CRS-Diff. CRS-Diff exhibits relatively limited image quality due to its constrained training data. The restricted diversity and quantity of training samples negatively impact the model’s ability to generalize across various geographic regions and land cover types. As a result, the generated images often lack the visual fidelity and semantic richness observed in outputs from models trained on more comprehensive datasets.
GeoSynth. GeoSynth is tailored explicitly for generating high-resolution remote sensing images, with training data predominantly sourced from urban areas in the United States. While it excels in producing detailed imagery within this domain, its generative capacity is limited when synthesizing scenes outside this geographic or semantic scope. In particular, its ability to represent diverse land use categories or non-urban environments is relatively weak, restricting its applicability in global or multi-domain remote sensing tasks.
Figure 13: Examples of generated object detection data. This includes the ability to generate data from different satellites.
Stable Diffusion. Although Stable Diffusion demonstrates strong general-purpose image generation capabilities, it lacks mechanisms for spatial layout control. This limitation is critical in remote sensing applications, where the accurate placement and arrangement of objects, such as buildings, roads, and vegetation, are essential for downstream tasks. The inability to control spatial semantics diminishes its utility in structured synthesis scenarios where geospatial coherence and object positioning matter.
ControlNet and EarthSynth. ControlNet and the proposed EarthSynth model support explicit spatial layout control, allowing for the guided generation of images with well-defined structures and localized semantic targets. This capability is particularly valuable in tasks such as data augmentation, simulation-based training, or synthetic dataset creation for segmentation and detection models. EarthSynth, in particular, further enhances visual realism while preserving layout fidelity, making it a powerful tool for generating structured, high-quality remote sensing imagery across diverse environments and object categories.
A.5.3 Guidance Scale Analysis
Figure 14: Comparison of different generation methods.
Figure 15 shows how the guidance scale affects the CLIP score. As the scale increases, the CLIP Score rises first, then levels off or slightly decreases. This suggests that moderate guidance scales lead to better alignment between the generated image and the text prompt. In the top examples, representing tennis courts, the CLIP Score reaches its peak around scale 4, with images showing more transparent structure and improved object fidelity. In the bottom examples, representing playgrounds, the score is highest near scale 2 or 3, but they have a poor image generated. Lower scales produce blurry or semantically weak images, while higher scales enhance visual clarity but may reduce diversity. These results indicate that a moderate guidance scale, typically between 3 and 5, balances semantic alignment and image quality well. And we can also find that generation varies across different images, and tuning the guidance scale provides a simple way to control semantic accuracy and visual structure.
A.6 Limitations
We summarize key considerations and limitations of the EarthSynth-180K and EarthSynth as follows:
Limited Multispectral Generalization. This work focuses on generating optical images. However, optical images only cover the visible spectrum and lack the broader spectral information in multispectral images. This limits their use in vegetation monitoring, material classification, and environmental analysis. Although the EarthSynth framework aims for cross-satellite generalization, training and testing are done on the EarthSynth-180K dataset without standard multispectral satellites. Therefore, generalization across sensors with different spectral features is unproven. Extending the framework to handle multispectral data is an important future direction.
More Training Cost. EarthSynth’s CF-Comp strategy assumes equal training costs, but training large generative diffusion models is time- and resource-intensive. Since the CF-Comp strategy involves sample combinations in each batch, these setups may not be feasible in resource-limited or costly environments, making some comparisons more theoretical.
Figure 15: The guidance scale affects the CLIP score across two categories.